Find the Re (w)where w is the positive solution to 22 =3+6i
Answer #1

given that 9 z 34 bi As we know that at ibarLO Where re 92 +62 b ton 3 tb √32 +62 Lein G - 145/tenl o 2 - 5452 ting As we kn2. Zo = zi ñ 1 sin (Otar has lon) + i sin Mees to +27) + 2 = 545(es ten ot i sin (ten (12) (45)2x0s ( terte pecake i sin (tenZo - 45 cas les .) + i sin sales = 45a tas tits) . Read (W) (45) x J5 b XL 45 2x5 - kasnesja (1125) 10 pini 4. lo =,=(4) 4 [